# Inflationary Gravitational Wave Spectral Shapes as test for Low-Scale Leptogenesis

### v1

@@ IV.1 Analytical Estimates of Equilibrium Temperature
Equilibrium temperature is the temperature at which matter domination ends, kickstarting the latest radiation dominated era. From the discussion above and from Fig. 3 and Fig. 4, we see that it is around the decay temperatures $T_{\phi}$ or $T_{N_{1}}$ depending on which one is smaller,
$T_{\text{dec.}}\sim\text{Min}[T_{\phi},T_{N_{1}}]$ (34)
We use this temperature in Eq. (48) for calculating the damped GW spectrum from inflation. Note that the exact equilibrium temperature is a little lower than the decay temperature since the decay is not instantaneous. However for our considerations, $T_{\rm dec.}$ is a reasonable estimate of it.
If we have thermal RHN scenario discussed in Case (c) above, provided the RHNs dominate before decaying, there are two stages of entropy injection, one from scalar decay at $T_{\phi}$ and another from RHN decay at $T_{N_{1}}$. In that case, we have an additional radiation dominated era between the two matter dominated epochs. Hence we take the two equilibrium temperatures to be $T_{\phi}$ and $T_{N_{1}}$ in this scenario.
@@ IV.2 Dilution Factor from Entropy Injection
Both scalar and RHN decay injects entropy into the Universe if the decay products thermalize quickly with the radiation bath. The general formula for the dimensionless dilution factor from the entropy injection by a dominating species $\chi$ decaying to relativistic particles is given by  Scherrer:1984fd ; Berbig:2023yyy
@@
For our analysis, we estimate the dilution factor using the formula in (35) for the 3 classifications we made:
Case (a): In this case, the RHNs decay out-of-equilibrium instantaneously to relativistic particles at $T_{\phi}$. Therefore we can assume that it is the scalar which is directly decaying to the radiation bath releasing entropy. The dilution factor $\Delta$ can be given by Eq. (35) with $\chi$ replaced by $\phi$. The initial abundance is given by Eq. (23). In the large entropy injection approximation $\Delta\gg 1$, using Eq. (27) we can write the dilution factor to be,
@@
Figure 5: Suppressed gravitational spectra of a benchmark Case (a) scenario when RHNs decay instantaneously. Here we have taken $r=0.035$, $n_{T}=0.5$, $T_{\rm RH}=10^{10}$ GeV, $K=10^{3}$, and varied $T_{\phi}$ and $M_{\phi}$. Gray line shows the standard GW spectrum without any intermediate matter domination. Black solid and dashed lines correspond to $M_{\phi}=10^{9}$ and $3\times 10^{9}$ GeV respectively. The extreme dampening at frequencies above $\sim 10$ Hz is due to inflaton decay.
Case (b): RHNs dominate the energy budget of the Universe as soon as they are produced. The energy density stored in $\phi$ species is first transferred to the RHNs and then dumped into the radiation bath when RHNs decay at $T_{N_{1}}$. Hence, this scenario can be imagined as if $\phi$ itself is decaying to SM particles at $T_{N_{1}}$.
The dilution factor in that case will be same as Case (a), except $\Gamma_{\phi}$ will be replaced by $\Gamma_{N_{1}}^{rf}$ in Eq. (35) to specify the elongated period of matter domination compared to Case (a). Note that considering only the lightest RHN $N_{1}$ is sufficient here. Again in the large entropy injection limit $\Delta\gg 1$, one can write the dilution factor as,
$\Delta\simeq 3.7\times 10^{9}\left(\frac{M_{\phi}}{10^{15}\text{ GeV}}\right)\left(\frac{10^{5}\text{ GeV}}{M_{1}}\right)\sqrt{\frac{10^{-4}}{K}}$ (37)
Figure 6: Same as Fig. 5 but for $n_{T}=0$. Solid and dashed black lines correspond to $M_{\phi}=10^{9}$ and $3\times 10^{9}$ GeV respectively.
Case (c): As discussed earlier, within suitable parameter choices, one can obtain thermal leptogenesis with two-step entropy injection (Appendix B). The dilution factor from scalar decay $\Delta_{\phi}$ is again estimated by Eq. (36) as in Case (a), while the dilution factor for RHNs, $\Delta_{N}$ is calculated assuming an initial RHN abundance just before their decay to be Bezrukov:2009th ,
$\frac{n_{N}}{s}=\frac{135\,\zeta(3)}{4\pi^{4}\,g_{\ast S}}$ (38)
which is based on the assumption that RHNs were relativistic when they froze out, analogous to Eq. (23). In the large entropy injection limit $\Delta_{N}$ is given as,
$\Delta_{N}\simeq 1109\,\sqrt{\frac{10^{-10}}{K}}$ (39)
Lastly, if the RHNs do not dominate (see Eq. (32)), we only have 1-step entropy injection from scalar decay and corresponding dilution factor $\Delta_{\phi}$. In that case, GW experiments can not probe the leptogenesis scale, since no suppression of the inflationary GW spectrum occurs at $T_{N_{1}}$.
@@ V.1 Primordial spectrum in standard and non-standard cosmology
Figure 7: Same as Fig. 5 but for a benchmark Case (b), “RHN domination” scenario. Here we take $T_{\phi}=10^{7}$ GeV, $r=0.035$, $n_{T}=0.5$ and vary the washout parameter $K$ along with $M_{1}$ and $M_{\phi}$. The spectral shape depends on the leptogenesis scale $T_{N_{1}}=M_{1}\sqrt{K}$ as shown. Again, solid and dashed black lines represent $M_{\phi}=10^{9}$ and $3\times 10^{9}$ GeV respectively.
@@
$\Omega_{GW}(k)=\frac{1}{12}\left(\frac{k}{a_{0}H_{0}}\right)^{2}T_{T}^{2}(k)P_{T}^{\text{prim.}}(k)$ (40)
with $a_{0}=1$ representing the present-day scale factor and $H_{0}\approx 2.2\times 10^{-4}$ Mpc-1 being the current Hubble expansion rate Datta:2022tab . $P_{T}^{\text{prim.}}(k)$ is the primordial tensor power spectrum from inflation, and $T_{T}^{2}(k)$ is the transfer function that characterizes the propagation of GWs in the background of a Friedmann-Lemaitre-Robertson-Walker (FLRW) Universe after the temperature at the time of horizon re-entry, $T_{\text{in}}$ given by Nakayama:2008wy :
@@
where $A_{S}(k_{*})$ is scalar power spectrum Planck:2018jri . For our analysis we fix the value of $r=0.035$. The spectral index $n_{T}$ plays a critical role in characterizing the inflationary tensor power spectrum. While standard single-field slow-roll inflation predicts a red-tilted spectrum, where $n_{T}$ satisfies the consistency relation $n_{T}=-r/8$ Liddle:1993fq , other scenarios, including blue-tilted spectra ($n_{T}>0$), are well-motivated in various cosmological models Brandenberger:2006xi ; Baldi:2005gk ; Kobayashi:2010cm ; Calcagni:2004as ; Calcagni:2013lya ; Cook:2011hg ; Mukohyama:2014gba ; Kuroyanagi:2020sfw .
Figure 8: Same as Fig. 5 but for Case (c), thermal leptogenesis, portraying 2-step entropy injection and a knee-like feature around $f\sim 10^{-2}$ Hz due to $\phi$ decay, in addition to the one due to RHN decay at lower frequencies. Here, we take $r=0.035$, $n_{T}=0.5$, $M_{\phi}=10^{13}$ GeV, $T_{\phi}=10^{5}$ GeV and $R_{\rm RH}=10^{15}$ GeV. The effect of $K$ is shown. The solid and dashed line correspond to $M_{1}=10^{5}$ and $3\times 10^{5}$ GeV.
Numerical computation and analytical fitting gives the transfer function $T_{T}^{2}(k)$ as:
@@
Here, $g_{*}^{0}=3.36$ and $g_{*S}^{0}=3.91$ are the present-day values of the effective degrees of freedom. $\Omega_{m}=0.31$ Datta:2022tab is the total matter density. The Bessel function $j_{1}(z_{k})$ describes the damping of GW amplitudes after horizon re-entry with $z_{k}\equiv k\tau_{0}$ and $\tau_{0}=2/H_{0}$. In the limit $z_{k}\gg 1$, relevant for the frequencies of interest, we can replace $j_{1}(z_{k})$ for $1/(\sqrt{2}z_{k})$. Finally, the fitting function $F(k)$ in standard cosmology is given by Kuroyanagi:2014nba ,
@@
Figure 9: Same as Fig. 8 but here we take $M_{1}=10^{6}$ GeV, $K=10^{-8}$ and show the dependence of the spectral shape on $T_{\phi}$ and $M_{\phi}$. The solid and dashed lines represent $M_{\phi}=10^{13}$ and $3\times 10^{13}$ GeV respectively.
@@ V.2 Two-step entropy injection
Our Case (c) discussed in Sec. IV has two stages of entropy injection. We propose a modified transfer function to incorporate the two stages along the lines of Kuroyanagi:2014nba , by assuming the form of the fitting function to be,
@@
where $k_{\text{dec.}}^{\phi},k_{\text{dec.}}^{N}$ are related to the decay temperatures $T_{\phi}$ and $T_{N_{1}}$ respectively and $k_{\text{dec.S}}^{\phi},k_{\text{dec.S}}^{N}$ are related to the domination temperatures of $\phi$ and $N_{1}$ respectively. We define them as,
@@
where $\Delta_{\phi},\Delta_{N}$ are the dilution factors from $\phi$ and RHN decays respectively. The necessary conditions to use this transfer function are $M_{\phi}\gg M_{1},T_{\phi}>T_{N_{1}}$ and Eq. (32) as discussed earlier. A proper way to obtain the 2-step transfer function requires numerical fitting which we left as a future direction. In this case we have two suppression frequencies corresponding to scalar and RHN decays, analogous to Eq. (55),
@@
Figure 10: Projected sensitivities of different experiments in the RHN parameter space for $n_{T}=0$ (Left) and $n_{T}=0.5$ (Right). Here $\tilde{m}_{1}$ is the effective neutrino mass defined by Eq. (8) and $m_{*}=1.1\times 10^{-3}$ eV is the equilibrium neutrino mass defined in Eq. (9). For both plots we took $M_{\phi}$= $10^{15}$ GeV , $T_{\phi}$ = $10^{10}$ GeV, $r=0.035$. The colored regions associated with each experiment implies sensitivity with SNR $>10$. As expected, a higher $n_{T}$ increases overall SNR. See main text for detailed description and analysis.
@@ V.3 Signal-to-noise ratio (SNR)
Interferometers are instruments that measure displacements in terms of dimensionless strain-noise denoted as $h_{\text{GW}}(f)$. This strain-noise is related to the amplitude of gravitational waves (GWs) and can be converted into an energy density, expressed by the formula:
$\Omega_{\text{exp}}(f)h^{2}=\frac{2\pi^{2}f^{2}}{3H_{0}^{2}}h_{\text{GW}}(f)^{2}h^{2}$ (65)
Here, $H_{0}$ represents the present-day Hubble rate ($H_{0}=h\times 100\frac{\text{km/s}}{\text{Mpc}}$). To assess the likelihood of detecting the primordial GW background, we calculate the signal-to-noise ratio (SNR) using the experimental sensitivity $\Omega_{\text{exp}}(f)h^{2}$, either given or projected. The SNR is determined by the following formula:
[…]
In this analysis, we use $h=0.7$ and an observation time of $\tau=4$ years. A detection threshold of SNR $\geq 10$ is applied.
In the spectrum plots we display the sensitivity curves for different ongoing and future GW experiments. They can be grouped as:
The blue violine lines in the spectrum plots are the most recent data from NANOGrav.

### v2

@@ IV.1 Analytical Estimates of Equilibrium Temperature
Equilibrium temperature is the temperature at which a matter-dominated epoch ends, kickstarting a radiation-dominated era. From the discussion above, we see that for Case (a) and (b), the equilibrium temperature is around the decay temperatures $T_{\phi}$. The exact equilibrium temperature is a little lower than the decay temperature since the decay is not instantaneous. However, for our considerations, $T_{\phi}$ is a reasonable estimate of the equilibrium temperature.
In Cases (c), there are two matter-dominated epochs, hence entropy is injected into the Universe in two steps as discussed earlier. First, the majoron decaying to relativistic RHNs increases the total entropy of the Universe at around $T_{\phi}$. The second stage of entropy injection is when RHNs decay at temperature $T_{N_{1}}$. There is a radiation-dominated epoch between the two matter-dominated epochs. Hence we take the two equilibrium temperatures to be $T_{\phi}$ and $T_{N_{1}}$. Thus we approximate the equilibrium temperatures for GW analysis as,
$T_{\rm eq.}\equiv T_{\text{dec.}}\sim\begin{cases}T_{\phi}\quad\text{for Case (a) and (b)}\\ T_{\phi}\,\&\,T_{N_{1}}\quad\text{for Case (c)}\end{cases}$ (48)
@@ IV.2 Dilution Factor from Entropy Injection
Both majoron and non-relativistic RHN decay inject entropy into the Universe if the decay products are relativistic. The general formula for the dimensionless dilution factor from the entropy injection by a dominating species $\chi$ decaying to relativistic particles is given by  Scherrer:1984fd ; Berbig:2023yyy
@@
For our analysis, we estimate the dilution factor using the formula in (49) for the four classifications we made:
Case (a): In this case, the RHNs decay out-of-equilibrium instantaneously to relativistic particles at $T_{\phi}$. Therefore we can assume that it is the majoron that is directly decaying to the radiation bath releasing entropy. The dilution factor $\Delta$ can be given by Eq. (49) with $\chi$ replaced by $\phi$. The initial abundance is given by Eq. (32). In the large entropy injection approximation $\Delta\gg 1$, using Eq. (35) we can write the dilution factor to be,
@@
Case (b): RHNs dominate the energy budget of the Universe as radiation after their production. Eventually, the RHNs decay relativistically at $T_{N}^{\rm rel}<T_{\phi}$. Hence the RHN decay injects negligible entropy into the Universe. The only stage of substantial entropy injection is when the majoron $\phi$ decays at $T_{\phi}$. The dilution factor is again estimated by Eq. (50).
Case (c): RHNs dominate the energy budget of the Universe as soon as they are produced, first as radiation, then as matter. The energy density stored in $\phi$ species is first transferred to the RHNs and then dumped into the radiation bath when RHNs decay at $T_{N_{1}}$. Hence, this scenario leads to two matter-dominated epochs separated by a radiation-dominated epoch.
The dilution factor from $\phi$ decay is again given by Eq. (50). The dilution factor from RHN decay can be estimated using Eq. (49) considering the RHN decay width $\Gamma_{N}^{rf}$ and initial abundance after $\phi$ decay as,
$\frac{n_{N}}{s_{\rm tot}^{\text{after}}}=\frac{2\,n_{\phi}}{\Delta_{\phi}\,s_{\rm tot}^{\rm before}}=\frac{2}{\Delta_{\phi}}\frac{n_{\phi}}{s}=\frac{45}{\pi^{4}}\frac{\zeta(3)}{g_{*S}}\frac{1}{\Delta_{\phi}}$ (51)
where $s_{\rm tot}$ is the total entropy of the Universe including entropy of the decoupled RHN sector, and superscripts “before” and “after” refers to before and after $\phi$ decay. Total entropy after $\phi$ decay, $s_{\rm tot}^{\rm after}$ can be written as $s_{\rm tot}^{\rm before}\Delta_{\phi}$ where $\Delta_{\phi}$ is given by Eq. (50). For the last equality, we use Eq. (32) since before the majoron decays the total entropy of the Universe is simply the entropy of the thermal bath. Note that considering only the lightest RHN $N_{1}$ is sufficient here. Again in the large entropy injection limit $\Delta_{N}\gg 1$, one can write the dilution factor from RHN decay as,
$\Delta_{N}\simeq 2000\,\left(\frac{10^{-12}}{K}\right)^{1/2}\left(\frac{10^{8}\text{ GeV}}{M_{\phi}}\right)\left(\frac{T_{\phi}}{100\text{ TeV}}\right)$ (52)
@@ V.1 Primordial spectrum in standard and non-standard cosmology
$\Omega_{GW}(k)=\frac{1}{12}\left(\frac{k}{a_{0}\mathcal{H}_{0}}\right)^{2}T_{T}^{2}(k)P_{T}^{\text{prim.}}(k)$ (53)
with $a_{0}=1$ representing the present-day scale factor and $\mathcal{H}_{0}\approx 2.2\times 10^{-4}$ Mpc-1 being the current Hubble expansion rate Datta:2022tab . $P_{T}^{\text{prim.}}(k)$ is the primordial tensor power spectrum from inflation, and $T_{T}^{2}(k)$ is the transfer function that characterizes the propagation of GWs in the background of a Friedmann-Lemaitre-Robertson-Walker (FLRW) Universe after the temperature at the time of horizon re-entry, $T_{\text{in}}$ given by Nakayama:2008wy :
@@
where $A_{S}(k_{*})$ is scalar power spectrum Planck:2018jri . For our analysis, we fix the value of $r=0.035$. The spectral index $n_{T}$ plays a critical role in characterizing the inflationary tensor power spectrum. While standard single-field slow-roll inflation predicts a red-tilted spectrum, where $n_{T}$ satisfies the consistency relation $n_{T}=-r/8$ Liddle:1993fq , other scenarios, including blue-tilted spectra ($n_{T}>0$), are well-motivated in various cosmological models Brandenberger:2006xi ; Baldi:2005gk ; Kobayashi:2010cm ; Calcagni:2004as ; Calcagni:2013lya ; Cook:2011hg ; Mukohyama:2014gba ; Kuroyanagi:2020sfw .
Numerical computation and analytical fitting give the transfer function $T_{T}^{2}(k)$ as:
@@
Here, $g_{*}^{0}=3.36$ and $g_{*S}^{0}=3.91$ are the present-day values of the effective degrees of freedom. $\Omega_{m}=0.31$ Datta:2022tab is the total matter density. The Bessel function $j_{1}(z_{k})$ describes the damping of GW amplitudes after horizon re-entry with $z_{k}\equiv k\tau_{0}$ and $\tau_{0}=2/\mathcal{H}_{0}$. In the limit $z_{k}\gg 1$, relevant for the frequencies of interest, we can replace $j_{1}(z_{k})$ for $1/(\sqrt{2}z_{k})$. Finally, the fitting function $F(k)$ in standard cosmology is given by Kuroyanagi:2014nba ,
@@ V.2 Two-step entropy injection
Our Case (c) discussed in Sec. IV have two stages of entropy injection. We could not find any previous study of the transfer function to fit two stages of entropy injection. Therefore, extrapolating on results obtained from numerical fitting in Kuroyanagi:2014nba , we propose a modified transfer function to incorporate the two stages by assuming the form of the fitting function analogous to Eq. (59) to be,
@@
where $k_{\text{dec.}}^{\phi},k_{\text{dec.}}^{N}$ are related to the decay temperatures $T_{\phi}$ and $T_{N_{1}}$ respectively and $k_{\text{dec.S}}^{\phi},k_{\text{dec.S}}^{N}$ are related to the domination temperatures of $\phi$ and $N_{1}$ respectively. Note that in Eq. (70), we use the function $T^{2}_{2}(x)$ for the energy scales corresponding to decay temperatures and $T^{2}_{3}(x)$ for energy scales corresponding to domination temperatures, as suggested in Kuroyanagi:2014nba . We define the energy scales analogous to Eq.s (61)-(64),
@@
where $\Delta_{\phi},\Delta_{N}$ are the dilution factors from $\phi$ and RHN decays respectively. Here in Eq. (75), we utilized the fact that the total dilution factor after $\phi$ and RHN decay is the product of the individual dilution factors $\Delta_{\phi}$ and $\Delta_{N}$ corresponding to the two stages of entropy injection. A proper way to obtain the 2-step transfer function requires numerical fitting which we left as a future direction. In this case, we have two suppression frequencies corresponding to majoron and RHN decays, analogous to Eq. (68),
@@ V.3 Signal-to-noise ratio (SNR)
Interferometers are instruments that measure displacements in terms of dimensionless strain-noise denoted as $h_{\text{GW}}(f)$. This strain-noise is related to the amplitude of gravitational waves (GWs) and can be converted into an energy density, expressed by the formula,
$\Omega_{\text{exp}}(f)h^{2}=\frac{2\pi^{2}f^{2}}{3\mathcal{H}_{0}^{2}}h_{\text{GW}}(f)^{2}h^{2}$ (78)
Here, $\mathcal{H}_{0}$ represents the present-day Hubble rate ($\mathcal{H}_{0}=h\times 100\frac{\text{km/s}}{\text{Mpc}}$). To assess the likelihood of detecting the primordial GW background, we calculate the signal-to-noise ratio (SNR) using the experimental sensitivity $\Omega_{\text{exp}}(f)h^{2}$, either given or projected. The SNR is determined by the following formula,
$\text{SNR}\equiv\sqrt{\tau\int_{f_{\text{min}}}^{f_{\text{max}}}df\left(\frac{\Omega_{\text{GW}}(f)h^{2}}{\Omega_{\text{exp}}(f)h^{2}}\right)^{2}}$ (79)
In this analysis, we use $h=0.7$ and an observation time of $\tau=4$ years. A detection threshold of SNR $\geq 10$ is applied. It is important to note that this formula for SNR calculation is derived under the weak signal approximation, where the GW signal is much smaller than the instrumental noise Allen:1997ad . While it may overestimate the true SNR for strong signals, we adopt it here for both weak and strong GW signals for simplicity.
In the spectrum plots, we display the sensitivity curves for different ongoing and future GW experiments. They can be grouped as:
The blue violin lines in the spectrum plots are the most recent data from NANOGrav.
@@ V.4 Bounds from BBN and CMB on dark radiation
Figure 8: Suppressed gravitational spectra of a benchmark Case (a) scenario when RHNs decay instantaneously or Case (b) where RHNs decay relativistically. Here we have taken $r=0.035$, $T_{\rm RH}=10^{13}$ GeV, $M_{1}=10^{5}$ GeV, $K=10^{3}$, and varied $T_{\phi}$ and $M_{\phi}$. Value of spectral index is $n_{T}=0$ (Left) and $n_{T}=0.3$ (Right). The gray line shows the standard GW spectrum without any intermediate matter domination. Black solid and dashed lines correspond to $M_{\phi}=10^{9}$ and $3\times 10^{9}$ GeV respectively. The extreme dampening at frequencies above $\sim 100$ Hz is due to inflaton decay.
Figure 9: Same as Fig. 8 but for a benchmark Case (c), “RHN matter domination” scenario. The Left plot shows the dependence of the spectra on $M_{1}$ and $K$ while the Right plot shows the dependence on $M_{\phi}$ and $T_{\phi}$. In the Left plot, the solid and dashed black lines represent $M_{1}=10^{8}$ and $3\times 10^{8}$ GeV respectively. In the Right plot, the solid and dashed black lines represent $M_{\phi}=10^{12}$ and $2\times 10^{12}$ GeV respectively.